OpenStudy (anonymous):

Simplify the expression. sqrt6(5 - sqrt 6) + 4 sqrt6

OpenStudy (cwrw238):

= 5sqrt6 - 6 + 4 sqrt6 can you continue?

OpenStudy (cwrw238):

treat the terms in sqrt6 as if they were terms in x eg 2x +3x = 5x

OpenStudy (cwrw238):

5x + 4x = 9x in a similar fashion: 5sqrt6 + 4 sqrt6 = 9 sqrt6

OpenStudy (anonymous):

so I believe the answer would be 9 sqrt 6-6, did I do that correctly?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Thank you so much for helping me work through this problem...

OpenStudy (cwrw238):

dod you understand where the -6 comes from?

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

I do

OpenStudy (anonymous):

\[(\sqrt{6})(5-\sqrt{6})+4\sqrt{6}\]

OpenStudy (anonymous):

=\[5\sqrt{6}-\sqrt{6}^{2}+4\sqrt{6}\]

OpenStudy (anonymous):

\[5\sqrt{6}+4\sqrt{6}\]

OpenStudy (anonymous):

oops sorry +6

OpenStudy (anonymous):

\[9\sqrt{6}+6\]

OpenStudy (anonymous):

no sorry it's -6 at the end. so that would be your answer

OpenStudy (bradely):

step by step answer is in below http://www.mathskey.com/question2answer/11535/simplify-the-expression